Find the equation of the tangent line to a curve y=-x2-1 that is parallel to the line 2x+y=6.


1
Expert's answer
2022-03-16T10:12:19-0400

equation of the tangent of the curve at the point a can be found the following way

"f(x)-f(a)=f'(a)(x-a)" , f'(a) is the slope of this tangent line, so, since this tangent line is parallel to the line 2x + y = 6, then f'(a) = -2

"y'=-2x\implies f'(a)=-2a\implies -2a=-2\implies a=1"

"f(1)=-2"

so, the sought equation is "y+2=-2(x-1)\implies y=-2x"